# Shaving Weights with Occam’s Razor: Bayesian Sparsification for Neural Networks using the Marginal Likelihood

### v1

@@ 3.2 Learning Regularization with the Marginal Likelihood
Figure 2: Predictive performance as a function of sparsity level in unstructured pruning. We see that SpaM improves the performance over MAP training across most architectures and datasets and that OPD often outperforms the other pruning criteria. Both of these effects are particularly visible at higher sparsity levels. We also see that our online pruning approach is often competitive with post-hoc pruning.
To optimize the potentially millions of regularization parameters, for example, arising from a parameter-wise prior, we employ the marginal likelihood as a differentiable objective. Optimizing regularization parameters has the advantage that different (groups of) parameters will be regularized differently and therefore become easier to prune. While it would be intractable to optimize that many regularization parameters using validation-based forms of optimization, the marginal likelihood can be estimated and differentiated during training (Immer et al., 2021, 2023; Lin et al., 2023).
Automatically determining the relevance of parameter-groups (ARD) is a common approach in Bayesian learning that can lead to sparsity and smaller parameter magnitudes (MacKay et al., 1994; Tipping, 2001) and has been used especially in linear models. The Bayesian marginal likelihood provides an objective that automatically regularizes irrelevant parameter-groups more to lower their magnitude. Therefore, it is said to implement a Bayesian variant of Occam’s razor by finding the simplest model explaining the data well (Rasmussen and Ghahramani, 2000).
@@
or adaptive optimizers like Adam (Kingma and Ba, 2014). Algorithmically, we follow Immer et al. (2021) and optimize the Laplace approximation to the marginal likelihood after an initial burn-in phase with a certain frequency.
@@ 3.3 Optimal Posterior Damage (OPD)
While sparsity regularization learned by marginal likelihood training can advantageously be combined with any pruning criteria like Single-shot Network Pruning (SNIP; Lee et al., 2018), variants of Gradient Signal Preservation (GraSP; Wang et al., 2020; Lubana and Dick, 2021; Rachwan et al., 2022), or magnitude pruning (Han et al., 2016), we propose in this section a new pruning criterion which uses our Laplace approximation and extends the unstructured Optimal Brain Damage (OBD) pruning criterion (LeCun et al., 1990). While OBD traditionally uses the prior precision ${\mathbf{H}}_{{\boldsymbol{\theta}}}$, we propose to adapt it to use the posterior precision ${\mathbf{P}}_{{\boldsymbol{\theta}}}$. In this case the importance score $S(\theta_{p})$ for a given parameter becomes
@@
where ${\textnormal{P}}_{pp}$ denotes the posterior precision for the parameter $\theta_{p}$, extracted from the diagonal of the posterior precision matrix ${\mathbf{P}}_{{\boldsymbol{\theta}}}$. We call this novel posterior-based pruning criterion Optimal Posterior Damage (OPD). Intuitively, individual weights with high scores indicate certainty of the posterior distribution and a significant contribution to the model’s functionality, as indicated by the magnitude. In the unstructured process, we perform global one-shot pruning.
Further, we propose a structured version of OPD by aggregating the score of a structured set of parameters $g$, i.e.,
@@
In practice, the structured set of parameters $g$ corresponds to all parameters along one dimension of the weight matrix inside a layer, in order to reduce the size of the matrix multiplications. Since subsequent layers might have significantly different weight matrix dimensions impacting the magnitude of the aggregated sum, we opt for uniform structured pruning to guarantee a fair pruning treatment across all layers. Moreover, as removing a full structure is more aggressive, we also apply gradual pruning during training. Finally, we omit pruning the final layer to mitigate overly strong impacts on classification accuracy and computational stability (Elsen et al., 2019).
The OPD pruning criterion can be computed to prune online during training or post-hoc after training by using Laplace approximations. On one hand, in the case of maximum a posteriori (MAP) training, Laplace approximations of the inverse Hessian at ${\boldsymbol{\theta}}_{\ast}$ can be additionally computed to approximate OPD. On the other hand, in the case of our marginal likelihood training, Laplace approximations of the precision matrix can be reused to compute OPD without computational overhead, in contrast to the other pruning criteria, which often require additional computations to be performed. Note that we will also show in our experiments that OPD additionally avoids the need for potentially expensive fine-tuning after pruning.
@@ 4 Related work
Table 1: Accuracies of pruned ResNets on CIFAR-10. The best training method for each pruning criterion is highlighted in green, where we see that SpaM improves performance for all criteria except the random baseline. The best performances overall at each sparsity level are shown in bold, showing that our OPD criterion outperforms the others at high sparsities.
@@
From the early inception of Bayesian neural networks (Neal, 1992; Hinton and Van Camp, 1993), the Laplace approximation was a popular inference method (MacKay, 1992). In recent years, it has undergone a renaissance (Martens and Grosse, 2015; Botev et al., 2017; Ritter et al., 2018; Daxberger et al., 2021), including critical work on using more scalable approximations for the associated marginal likelihood in the context of model selection (MacKay, 1995; Immer et al., 2021, 2022), which we use in our framework. To the best of our knowledge, we are the first to study the benefits of this Laplace-approximated marginal likelihood in the context of sparsification of deep neural networks. However, similar methods that automatically quantify the relevance (ARD) of parameters have been derived and used for linear and kernel models (Tipping, 2001; Wipf and Rao, 2004).
@@
Various pruning criteria have been proposed to determine the importance of model parameters. Many criteria proposed to prune based on the weight magnitude (Han et al., 2016; Zhou et al., 2020; Bellec et al., 2018) but usually required additional fine-tuning to recover accuracy. Sun et al. (2023) proposed to combine activation and weight norms for pruning without fine-tuning. Other approaches include pruning using first-order information based on connectivity (Lee et al., 2018) or synaptic flow conservation (Tanaka et al., 2020), or second-order information aiming at preserving gradient flow (Rachwan et al., 2022; Wang et al., 2020; Lubana and Dick, 2021). Recently, van der Ouderaa et al. (2023) focused on pruning LLMs based on a second-order Taylor expansion. In contrast, OPD uses second-order information provided by the posterior precision given by the Laplace approximation.
Beyond pruning criteria, there have been many approaches to prune at initialization (Lee et al., 2018; Wang et al., 2020; Tanaka et al., 2020), during training (Golkar et al., 2019; Zhou et al., 2021), and after training (Han et al., 2016; Sun et al., 2023). In particular, multiple works proposed to leverage specific training schemes promoting zero-invariant parameter groups for structured pruning (Chen et al., 2021, 2023). In contrast, SpaM induces sparse structures during training using Bayesian marginal likelihood training.
@@ 5 Experiments
We conduct experiments on various datasets and models and outline our experimental setup in detail in Appendix D. We compare MAP training with our proposed SpaM approach with different priors, comparing our OPD pruning criterion with random pruning, magnitude pruning, SNIP (Lee et al., 2018), GraSP (Wang et al., 2020; Lubana and Dick, 2021), and SynFlow (Tanaka et al., 2020). We show that SpaM improves pruning performance across a range of different pruning criteria, especially at higher sparsities, and that our OPD often outperforms the baselines. This observation extends not only to predictive accuracy, but also uncertainty estimation. Moreover, we show that the choice of prior can play a significant role and introduce parameter-wise and unit-wise priors for the KFAC approximation. Finally, we show that SpaM and OPD also work in a structured pruning setting, leading to significant computational benefits.
@@ 5.1 SpaM Improves Performance at High Sparsities
Figure 3: Uncertainty estimation with pruned ResNets on CIFAR-10. We see that SpaM improves uncertainty estimation in terms of NLL, ECE, and Brier score for many pruning criteria and that our OPD criterion outperforms the baselines, especially at high sparsities.
@@
Figure 4: Comparison of different priors and Hessian approximations for SpaM-OPD pruning. The unit-wise and parameter-wise priors show better performance at high sparsity levels, with the parameter-wise one bridging the gap between Diag and KFAC LA.
@@
Figure 5: Similarly to unstructured pruning, we also see in this experiment on structured pruning that SpaM (using a unit-wise prior) improves performance over MAP and that OPD mostly outperforms other pruning criteria, especially at higher sparsity levels.
@@
Figure 6: Structured pruning with LeNet on FashionMNIST, using unit-wise priors. We see that our SpaM-OPD dominates the Pareto frontier in terms of predictive performance as a function of computational time and memory cost.
@@
We compare SpaM to MAP training with different pruning criteria, including OPD, across different models and datasets. For SpaM in this unstructured pruning context, we use the diagonal Laplace approximation with a parameter-wise prior. Encouragingly, MAP and SpaM reach comparable performance during training, showing that the increased sparsifiability of SpaM comes at no additional cost in unpruned performance (see Figure 7 in the appendix).
@@
More excitingly, we see in Figure 2 and Table 1 that SpaM drastically improves the performance for many pruning criteria, especially magnitude pruning, GraSP, and OPD. We also see that OPD, despite being a cheap byproduct of our marginal likelihood computation, often outperforms the other pruning criteria, especially at higher sparsities. For instance, at 95 % pruning rate (i.e., with 20x fewer parameters), our combination of SpaM and OPD still retains almost the same performance as the unpruned model, while the other pruning criteria with MAP training have dropped to essentially unusable performance levels at this sparsity.
@@
Fine-tuning. We see in Figure 11 in the appendix that some of this performance difference can be remedied by costly fine-tuning of the networks after pruning, which however still does not allow the other methods to reach the full SpaM-OPD performance, and in the case of OPD, also does not further improve its already near-optimal performance.
@@
Online pruning. Figure 2 shows that our online version of SpaM, which uses the marginal likelihood and OPD during training to iteratively prune the network, reaches comparable performance levels to the post-hoc version, thus offering a computationally even more convenient way to effectively sparsify neural networks.
@@
Uncertainty estimation. Given that SpaM is a Bayesian method, it does not only offer high predictive accuracies but also calibrated uncertainty estimates. Indeed, we see in Figure 3 that the trends we have seen for accuracy also apply for negative log-likelihood, expected calibration error, and the Brier score. Again, SpaM improves the uncertainty estimates over MAP training, OPD outperforms most other criteria, and we achieve well-calibrated models up until very high sparsity levels. Note that the random baseline also achieves a low ECE at high sparsity levels because it essentially reverts to random guessing, which is a known weakness of the ECE metric (Gruber and Buettner, 2022).
@@ 5.2 Influence of Priors on Sparsifiability
The Bayesian marginal likelihood, as employed in SpaM, strongly depends on the chosen prior. To understand the influence of prior and Hessian approximation on performance in our proposed SpaM-OPD approach, we compare diagonal and KFAC approximations with scalar, layer-wise, unit-wise, and parameter-wise priors. Note regarding the latter two, that in this work, we are the first to ever implement them for the KFAC approximation, thus contributing to the general framework of Laplace-approximated BNNs (Daxberger et al., 2021), independent of the pruning use case.
@@ 5.3 SpaM Extends to Structured Sparsification
Encouragingly, we see in Figure 5 that our findings from the unstructured case transfer qualitatively also to the structured case, with SpaM-OPD outperforming the baselines at high sparsities. Crucially, while the sparsity patterns generated by unstructured pruning are more difficult to translate into computational benefits, structured pruning directly leads to computational savings on standard GPUs (see also Figure 18 in the appendix). We see in Figure 6 that SpaM-OPD dominates the Pareto frontier of the tradeoff between performance and computational cost at high sparsities (i.e., low costs), yielding 10x–20x savings in FLOPS and memory consumption with only minimal deterioration in performance. This positions our proposed framework as a potentially promising approach for the deployment of AI models in resource-constrained environments.
6 Conclusion, Limitations, and Future Work
We have shown that the Bayesian marginal likelihood, with its associated Occam’s razor effect, can be used to select neural network models that are inherently more sparsifiable. Crucially, we have shown that this sparsifiability extends across different pruning criteria and enables large gains in performance and uncertainty estimation, especially at high sparsity levels. Conveniently, the computations needed for the marginal likelihood estimation using the Laplace approximation can be re-used to define a novel pruning criterion called OPD, which outperforms many existing (more expensive) criteria in our experiments. We have also presented guidelines for choosing priors within our framework and have shown that even in the challenging setting of structured pruning, our proposed SpaM approach can yield up to 20x savings in computational time and memory, with only small reductions in performance. Our work thus offers a promising path towards pruning large AI models at high sparsity levels for deployment on resource-constrained devices.
@@ 6 Conclusion, Limitations, and Future Work
Future Work.
Interesting avenues for future work include the exploration of more complex priors, the extension of our unit-wise priors and structured pruning approach to more complex neural network architectures, such as graph neural networks, and the empirical evaluation of our approach on large-scale foundation models, such as large language models (LLMs).

### v2

@@ 3.2 Learning Regularization with the Marginal Likelihood
To optimize the potentially millions of regularization parameters, for example, arising from a parameter-wise prior, we employ the marginal likelihood as a differentiable objective. Optimizing regularization parameters has the advantage that different (groups of) parameters will be regularized differently and, therefore, become easier to prune. While it would be intractable to optimize that many regularization parameters using validation-based forms of optimization, the marginal likelihood can be estimated and differentiated during training [12, 31, 32].
Automatically determining the relevance of parameter-groups (ARD) is a common approach in Bayesian learning that can lead to sparsity and smaller parameter magnitudes [33, 34] and has been used especially in linear models. The marginal likelihood provides an objective that automatically regularizes irrelevant parameter groups more to lower their magnitude. Therefore, it implements a Bayesian variant of Occam’s razor, finding the simplest model that explains the data well [13].
@@
or adaptive optimizers like Adam [35]. We follow Immer et al. [12] and optimize the Laplace approximation to the marginal likelihood after an initial burn-in phase with a certain frequency. We describe the optimization process and the related hyperparameters in Appendix D.4
@@ 3.3 Optimal Posterior Damage (OPD)
While sparsity regularization learned by marginal likelihood training can be advantageously combined with any pruning criterion, like Single-shot Network Pruning [SNIP; 36], variants of Gradient Signal Preservation [GraSP; 37, 38, 39], or magnitude pruning [7], we further propose a new pruning criterion that uses our Laplace approximation and extends the unstructured Optimal Brain Damage (OBD) pruning criterion [3]. While OBD traditionally uses the Hessian (approximation) ${\mathbf{H}}_{{\boldsymbol{\theta}}}$ of the loss, we propose to adapt it to use the posterior precision ${\mathbf{P}}_{{\boldsymbol{\theta}}}$, which additionally includes the prior precision ${\boldsymbol{\delta}}$. The importance score $S(\theta_{p})$ for parameter $\theta_{p}$ becomes
@@
where ${\textnormal{P}}_{pp}$ denotes the posterior precision for the parameter $\theta_{p}$, extracted from the diagonal of the posterior precision matrix ${\mathbf{P}}_{{\boldsymbol{\theta}}}$. We call this novel posterior-based pruning criterion Optimal Posterior Damage (OPD). Intuitively, individual weights with high scores indicate certainty of the posterior distribution and a significant contribution to the model’s functionality, as indicated by the magnitude.
We also propose a structured version of OPD by aggregating the score over a set of parameters $g$, i.e.,
@@
In practice, the structured set of parameters $g$ corresponds to all parameters along one dimension of the weight matrix inside a layer, in order to reduce the size of the matrix multiplications. Since subsequent layers might have significantly different weight matrix dimensions impacting the magnitude of the aggregated sum, we opt for uniform structured pruning to guarantee a fair pruning treatment across all layers. This means we prune each layer by the same target percentage, reducing the dimensions of each layer by the same proportion relative to the unpruned model. This approach contrasts with achieving a global target sparsity that varies across layers, which can make it more difficult to consistently compress and adjust the model’s size.
Moreover, as removing a full structure is more aggressive, we also apply gradual pruning during training. Finally, we omit pruning the final layer to mitigate overly strong impact on classification accuracy and computational stability [40].
When using our SpaM approach, the precomputed precision matrix from the Laplace approximation can be reused to compute OPD without computational overhead, in contrast to the other pruning criteria, which often require additional computations to be performed. Note that we will also show in our experiments that OPD additionally avoids the need for potentially expensive fine-tuning after pruning. Moreover, even in the case of maximum a posteriori (MAP) training, Laplace approximations of the inverse Hessian at ${\boldsymbol{\theta}}_{\ast}$ can be additionally computed to approximate OPD. Finally, the OPD criterion can not only be computed post-hoc after training, but even online during training.
@@ 4 Related work
Table 1: Accuracies of pruned ResNets on CIFAR-10. The best training method for each pruning criterion is highlighted in green, where we see that SpaM improves performance for all criteria except the random baseline. The best performances overall at each sparsity level are shown in bold, showing that our cheap OPD criterion outperforms the others at high sparsities.
@@
From the early inception of Bayesian neural networks [41, 42], the Laplace approximation was a popular inference method [16]. In recent years, it has undergone a renaissance [18, 19, 20, 14], including critical work on using more scalable approximations for the associated marginal likelihood in the context of model selection [11, 12, 43], which we use in our framework. To the best of our knowledge, we are the first to study the benefits of this Laplace-approximated marginal likelihood in the context of sparsification of deep neural networks. However, similar methods that automatically quantify the relevance (ARD) of parameters have been derived and used for linear, bilinear, and kernel models [34, 44, 45] as an alternative to the Lasso. More recently, van der Ouderaa et al. [46] and Bouchiat et al. [47] used the ARD mechanism in deep learning to select layers and features by regularization, respectively.
@@
Various pruning criteria have been proposed to determine the importance of model parameters. Many criteria prune based on the weight magnitude [7, 48, 49] but usually required additional fine-tuning to recover accuracy. Sun et al. [8] proposed to combine activation and weight norms for pruning without fine-tuning. Other approaches include pruning using first-order information based on connectivity [36] or synaptic flow conservation [50], or second-order information aiming at preserving gradient flow [37, 38, 39]. Recently, van der Ouderaa et al. [51] focused on pruning LLMs based on a second-order Taylor expansion. In contrast, OPD uses second-order information provided by the posterior precision given by the Laplace approximation. Beyond pruning criteria, there have been many approaches to prune at initialization [36, 37, 50], during training [52, 53], and after training [7, 8]. In particular, multiple works proposed to leverage specific training schemes promoting zero-invariant parameter groups for structured pruning [54, 55]. In contrast, SpaM induces sparsifiability during training, and is agnostic about the criterion.
@@ 5 Experiments
Figure 2: Predictive performance as a function of sparsity level in unstructured pruning. We see that SpaM improves the performance over MAP training across most architectures, datasets, and pruning criteria, and that OPD often outperforms the other pruning criteria. Both of these effects are particularly visible at higher sparsity levels. The black star in each subfigure denotes the performance of the unpruned models, which is often identical to the performance of models pruned at 20% sparsity.
We conduct experiments on various datasets and models and outline our experimental setup in detail in Appendix D. We compare MAP training with our proposed SpaM approach with different priors, comparing our OPD pruning criterion with random pruning, magnitude pruning, SNIP [36], GraSP [37, 38], and SynFlow [50]. We show that SpaM improves pruning performance with different pruning criteria, especially at higher sparsities, and that our OPD criterion often outperforms the other criteria. This observation extends not only to predictive accuracy, but also uncertainty estimation. Moreover, we show that the choice of prior can play a significant role, and we introduce parameter-wise and unit-wise priors for the KFAC approximation. Finally, we show that SpaM and OPD also work in a structured pruning setting, leading to significant computational benefits. The code for our methods and experiments can be found at https://github.com/fortuinlab/spam-pruning.
@@ 5.1 SpaM Improves Performance at High Sparsities
Figure 3: Uncertainty estimation with pruned ResNets on CIFAR-10. We see that SpaM improves uncertainty estimation in terms of NLL, ECE, and Brier score for many pruning criteria and that our OPD criterion outperforms the other criteria, especially at high sparsities.
@@
Figure 4: Comparison of different priors and Hessian approximations for SpaM-OPD unstructured pruning. The unit-wise and parameter-wise priors show better performance at high sparsity levels, with the parameter-wise one bridging the gap between Diag and KFAC LA.
@@
Figure 5: Similarly to unstructured pruning, we see in this experiment on structured pruning that SpaM (using a unit-wise prior) improves performance over MAP and that OPD mostly outperforms other pruning criteria, especially at higher sparsity levels. The black stars reflect the performance of the unpruned models.
@@
Figure 6: Structured pruning with LeNet on FashionMNIST, using unit-wise priors. We see that our SpaM-OPD dominates the Pareto frontier, in terms of predictive performance as a function of computational time and memory cost, and is particularly competitive at lower costs.
@@
We compare SpaM to MAP training with different pruning criteria, including OPD, across different models on tabular, vision, and language datasets. For SpaM in this unstructured pruning context, we use the diagonal Laplace approximation with a parameter-wise prior. Encouragingly, MAP and SpaM reach comparable performance during training, showing that the increased sparsifiability of SpaM comes at no additional cost in unpruned performance (see Figure B1 in the appendix).
@@
We see in Table 1 and Figure 2 that SpaM drastically improves the performance for many pruning criteria, especially magnitude pruning, GraSP, and OPD. We also see that OPD, despite being a cheap byproduct of our marginal likelihood computation, often outperforms the other pruning criteria, especially at higher sparsities. For instance, at 95 % pruning rate (i.e., with 20x fewer parameters), our combination of SpaM and OPD still retains almost the same performance as the unpruned model on vision tasks, while the other pruning criteria with MAP training have dropped to essentially unusable performance levels at this sparsity.
@@
Fine-tuning. We see in Figure B4 in the appendix that some of this performance difference can be remedied by costly fine-tuning of the networks after pruning, which however still does not allow the other methods to reach the full SpaM-OPD performance. Interestingly, in the case of OPD, this does not further improve its already near-optimal performance.
@@
Online pruning. Figure B5 in the appendix shows that our online version of SpaM, which uses the marginal likelihood and OPD during training to iteratively prune the network, reaches comparable performance levels to the post-hoc version, thus offering a computationally even more convenient way to effectively sparsify neural networks.
@@
Uncertainty estimation. Given that SpaM is a Bayesian method, it does not only offer high predictive accuracies but also calibrated uncertainty estimates. Indeed, we see in Figure 3 that the trends we have seen for accuracy also apply for negative log-likelihood, expected calibration error, and the Brier score. Again, SpaM improves the uncertainty estimates over MAP training, OPD outperforms most other criteria, and we achieve well-calibrated models up until very high sparsity levels. Note that the random baseline also achieves a low ECE at high sparsity levels because it essentially reverts to random guessing, which is a known weakness of the ECE metric [56].
@@ 5.2 Influence of Priors on Sparsifiability
To understand the influence of the prior and Hessian approximation on performance in our proposed SpaM-OPD approach, we compare diagonal and KFAC approximations with scalar, layer-wise, unit-wise, and parameter-wise priors. Note regarding the latter two, that in this work, we are the first to implement them for the KFAC approximation, thus contributing to the general framework of Laplace-approximated BNNs [14], independent of the pruning use case.
@@ 5.3 SpaM Extends to Structured Sparsification
Encouragingly, we see in Figure 5 that our findings from the unstructured case transfer qualitatively also to the structured case, with SpaM-OPD outperforming the baselines at high sparsities. Crucially, while the sparsity patterns generated by unstructured pruning are more difficult to translate into computational benefits, structured pruning directly leads to computational savings on standard GPUs (see also Figure B11 in the appendix). We see in Figure 6 that SpaM-OPD dominates the Pareto frontier of the tradeoff between performance and computational cost at high sparsities (i.e., low costs), yielding 10x–20x savings in FLOPS and memory consumption with only minimal deterioration in performance. This positions our proposed framework as a potentially promising approach for the deployment of AI models in resource-constrained environments.
6 Conclusion
We have shown that the Bayesian marginal likelihood, with its associated Occam’s razor effect, can be used during training to select neural network models that are inherently more sparsifiable. Crucially, we have shown that this sparsifiability extends across different pruning criteria and enables large gains in performance and uncertainty estimation, especially at high sparsity levels. Conveniently, the computations needed for the marginal likelihood estimation using the Laplace approximation can be re-used to define a novel pruning criterion called OPD, which outperforms many existing (more expensive) criteria in our experiments. We have also presented guidelines for choosing priors within our framework and have shown that even in the challenging setting of structured pruning, our proposed SpaM approach can yield up to 20x savings in computational time and memory, with only small reductions in performance. Our work thus offers a promising path towards pruning large AI models at high sparsity levels for deployment on resource-constrained devices.
@@ 6 Conclusion
Acknowledgments
AI acknowledges funding through a Max Planck ETH Center for Learning Systems (CLS) fellowship. VF was supported by a Branco Weiss Fellowship.